# 11footnotetext: Corresponding author

### v1

@@ Abstract
Learning general-purpose reasoning capabilities has long been a challenging problem in AI. Recent research in large language models (LLMs), such as DeepSeek-R1, has shown that reinforcement learning techniques like GRPO can enable pre-trained LLMs to develop reasoning capabilities using simple question-answer pairs. In this paper, we aim to train visual language models (VLMs) to perform reasoning on image data through reinforcement learning and visual question-answer pairs, without any explicit chain-of-thought (CoT) supervision. Our findings indicate that simply applying reinforcement learning to a VLM—by prompting the model to produce a reasoning chain before providing an answer—can lead the model to develop shortcuts from easy questions, thereby reducing its ability to generalize across unseen data distributions. We argue that the key to mitigating shortcut learning is to encourage the model to interpret images prior to reasoning. Therefore, we train the model to adhere to a caption-reason-answer output format: initially generating a detailed caption for an image, followed by constructing an extensive reasoning chain. When trained on 273K CoT-free visual question-answer pairs and using only reinforcement learning, our model, named Visionary-R1, outperforms strong multimodal models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, on multiple visual reasoning benchmarks.
@@ 1 Introduction
Reasoning is essential for enabling AI to tackle complex problems and make informed decisions in real-world applications. However, training AI models to reason is extremely challenging—primarily due to the lack of large-scale human-annotated reasoning data [1, 2, 3]. Recent advances in large language models (LLMs), such as DeepSeek-R1 [4], have demonstrated the potential to induce reasoning capabilities in LLMs via reinforcement learning and using only question-answer pairs, without explicit step-by-step supervision. Meanwhile, the computer vision community has begun exploring RL approaches for visual language models (VLMs), using methods like GRPO [5] to extend reasoning to multimodal settings [6, 7, 8, 9]. While these efforts are promising, existing visual reasoning models often rely on complex multi-stage training pipelines that are both computationally expensive and time-consuming. Moreover, these models heavily rely on labeled chain-of-thought reasoning data distilled from proprietary models like GPT-4o—limiting scalability and openness.
@@
To address the shortcut issue, we propose Visionary-R1, a reinforcement learning framework that enforces visual understanding before reasoning. The key idea is to train the model in a structured caption–reason–answer format, where it must first generate a detailed caption of the image before reasoning and answering. The captioning step ensures that the model does not just rely on superficial cues or patterns but engages in a deeper analysis of the image context, regardless of whether the question is easy or hard—this forces the model to adopt a consistent problem-solving approach, thus mitigating potential shortcuts and consequently making the reasoning capabilities more generalizable across different data distributions. To ensure the caption is informative, we impose auxiliary supervision on the caption tokens by using reinforcement learning from AI feedback [10]. This caption reward is combined with standard accuracy and format rewards during policy optimization. The resulting model produces longer, more meaningful reasoning tokens than the model learned with GRPO alone (see Fig. 1), leading to better generalization performance on unseen data (see Fig. 2).
To evaluate our approach, we compile a comprehensive dataset that aggregates 11 popular question-answer datasets, covering areas such as scene understanding, chart analysis, mathematical problem-solving, and document processing. In total, the training data consists of 272.6K CoT-free question-answer pairs. After training, Visionary-R1 is evaluated on several challenging visual reasoning benchmarks including MathVista [11], MathVision [12], MMStar [13], and MMBench [14]. The results show that Visionary-R1 outperforms strong proprietary models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, as well as the latest competitors based on supervised pre-training and reinforcement fine-tuning.
@@ 2 Related Work
Learning LLMs/VLMs that can reason have gained increasing attention from both academia and industry due to their ability to generate human-like, step-by-step reasoning, which is advantageous for tackling complex problems and delivering more interpretable answers [15, 16]. Supervised fine-tuning (SFT) is the most straightforward method to enhance a model’s reasoning capabilities, which relies on labeled data containing thinking processes. Since collecting human annotations is costly, existing work often resorts to using a pre-trained model like OpenAI’s GPT-4o to generate reasoning labels. For instance, LLaVA-CoT [17] utilizes GPT-4o to label 100K visual question-answer datasets with detailed chain-of-thought including summary, caption, and reasoning. However, the process of collecting CoT labels can be quite expensive, and the use of GPT-4o limits scalability while introducing a significant performance upper bound. Similarly, MMCR [18] also creates a 310k multi-turn reasoning dataset using GPT-4o. CoMCTS [19] introduces the Mulberry-260k dataset, which is specifically crafted to train tree-structure reasoning models. Compared to these models, our Visionary-R1 only uses simple question-answer pairs for training without any chain-of-thought supervision, yet it achieves stronger reasoning performance.
@@
Compared to SFT, reinforcement learning (RL) has recently been proved more effective in developing general-purpose reasoning capabilities as this paradigm has the potential to enable the model to explore reasoning in a broader language space and develop its own thinking processes [20]. Insight-V [21] presents a multi-agent system to select preference data from self-generated reasoning paths and optimizes the model based on a preference learning algorithm. R1-VL [22] designs step-wise rewards to improve reasoning accuracy and validity but relies on labeled data for SFT. RL has also been applied in Vision-R1 [23] and R1-Onevision [24], but only 10K samples are used in these models for RL training while the main focus is on SFT (that uses more than 200K samples). Our Visionary-R1 departs from the popular SFT-followed-by-RL pipeline and adopts a pure RL approach.
@@ 3.1 Motivation: The Shortcut Phenomenon in Visual Reasoning
While the GRPO [5] algorithm has been shown effective in improving the reasoning capabilities of language models, we observe a critical failure mode when transferring to visual reasoning tasks. This phenomenon manifests as a shortcut—GRPO often leads to degenerate behaviors where the model ignores the visual input and relies primarily on textual patterns from the question to generate an answer. As shown in Fig. 1, the model trained with GRPO can produce correct answers for simple questions during training—yet this is achieved without grounding in the image. This shortcut behavior can be particularly problematic in visual reasoning tasks, where the correct answer often depends on subtle image features such as embedded text, numerical values, object relationships, or chart patterns. Without forcing the model to attend to these visual signals, reinforcement learning alone encourages reward hacking: the model learns to exploit training distribution artifacts instead of learning general-purpose reasoning. To address this challenge, we propose a simple but effective modification: force the model to explicitly interpret the image before it begins reasoning. We operationalize this through a caption reward design (Section 3.2), which is then explicitly incorporated into the RL training objective (Section 3.3).
@@ 3.2 Visionary-R1: Grounding Reasoning via Captioning
While the format enforces structure, it does not guarantee that the caption is sufficiently detailed to support reasoning. To address this issue, we introduce a specialized caption reward $r_{c}\in\{0,1\}$ based on reinforcement learning from AI feedback [10]. Specifically, we feed the generated caption into an LLM, and ask it to answer the question based solely on the caption. In implementation, we use the LLM component of the policy model. If the answer is correct, the caption is deemed informative and rewarded; otherwise, it is penalized. This encourages the model to produce useful, visually grounded descriptions. The final reward for a sampled sequence $i$ is computed as:
@@ 3.3 Training Objective with Caption Reward
Group Relative Policy Optimization, known as GRPO, was originally developed in DeepSeekMath [5] for text-only reasoning tasks, and later adopted in DeepSeek-R1 [4]. GRPO simplifies the reinforcement learning paradigm by getting rid of the critic model. This is done by generating a group of responses for each sample and then computing the normalized reward within the group to determine an advantage value. To adapt it to visual reasoning, our method introduces two key differences. (1) First, as described in Section 3.2, we design a new reward structure by adding a caption reward that explicitly evaluates whether the model has interpreted the visual input, addressing the shortcut issue. (2) Second, we introduce a cosine-annealed KL penalty to stabilize training and encourage longer, more meaningful outputs—avoiding the limitations of a static KL coefficient in multimodal settings. We now detail our training objective and implementation.
@@
The KL divergence in Eq. 4 serves as a penalty term to prevent the model from straying too far from the baseline policy model, thereby stabilizing the training. It is non-trivial to determine the balancing weight for this term: using a large weight forces the model to stay within a close neighborhood of the baseline model and therefore impedes the model’s ability to engage in more in-depth thinking and generating long, detailed reasoning; on the other hand, using a small weight can lead to unstable training and potentially result in reward hacking [25]. To overcome this challenge, we propose dynamically annealing the KL penalty coefficient over time using cosine annealing, which uses a large coefficient during the early, unstable training phase and gradually reduces the value to allow the model to produce longer outputs in later stages. Specifically, we replace $\beta$ in Eq. 3 with $\hat{\beta}$, which is calculated as
@@ 4.1 Experimental Settings
We evaluate our approach on four widely-used visual reasoning benchmarks that cover various visual formats and question types: MathVista (Testmini) [11], MathVision [12], MMStar [13], and MMBench (en) [14]. MathVista encompasses a variety of reasoning types, including logical, algebraic, and scientific reasoning questions. MathVision focuses on mathematical visual reasoning tasks. MMstar aims to evaluate perception, math understanding, scientific and technology-related questions, and logical reasoning. MMBench is a comprehensive evaluation suite concerned with visual and mathematical reasoning.
@@
We adopt Qwen2.5-VL-3B [26] as the base model. This pre-trained model has strong visual understanding capabilities but has not undergone post-training for reasoning. For the group reward computation, we generate 8 output sequences (i.e., $n=8$ in Eq. 3) and the sampling temperature is set to 0.9 following the common practice. All parameters are optimized with a learning rate of $5\times 10^{-7}$. The caption reward’s balancing weight $\alpha$ is set to 0.1. The KL coefficient $\beta$ is set to 0.04. The model is trained for around 1,500 GPU hours on the 272.6K dataset with NVIDIA A800-80G GPUs.
Table 1: Comparison with state-of-the-arts on four challenging visual reasoning benchmarks. SFT and RL mean supervised fine-tuning and reinforcement learning, respectively. CoT means chain-of-thought, which is either self-generated or distilled from third-party models like GPT-4o. QA means that the model is learned with question-answer pairs only. Despite having only 3B parameters and using only QA data for training, Visionary-R1 beats strong commercial models like GPT-4o and Claude3.5-Sonnet. Note that indicates results borrowed from the Seed’s report [27].
Size Strategy Data MathVista MathVision MMStar MMBench
@@
GPT-4o [28] - - - 63.8 31.2 65.1 84.3
GPT-o1 [29] - - - 71.8 63.2 67.5 83.8
Claude3.5-Sonnet [30] - - - 67.7 37.9 65.1 82.6
Claude3.7-Sonnet [31] - - - 74.5 58.6 68.8 82.0
Gemini-1.5-Pro [32] - - - 63.9 19.2 59.1 73.9
Gemini-2.5-Pro [33] - - - 82.7 73.3 77.5 90.1
@@
Qwen2.5-VL [26] 3B - - 62.3 21.2 55.9 79.1
InternVL2.5 [34] 4B - - 60.5 20.9 58.3 81.1
MiniCPM-V2.6 [35] 8B - - 60.6 17.5 57.5 81.5
LLaMA3.2 [36] 11B - - 51.5 - 49.8 65.8
@@
Ovis [18] 4B SFT CoT 66.6 - 59.5 79.3
Mulberry [19] 7B SFT CoT 63.1 - 61.3 -
R1-Onevision [24] 7B SFT+RL CoT 64.1 29.9 - -
Insight-V [21] 7B SFT+RL CoT 59.9 - 61.5 82.3
R1-VL [22] 7B SFT+RL CoT 63.5 24.7 60 -
LLaVA-CoT [17] 11B SFT CoT 54.8 - 57.6 75
@@
Base Model 3B - - 61.5 19.1 52.4 82.1
SFT 3B SFT QA 54.6 7.0 61.9 80.7
GRPO 3B RL QA 61.8 20.3 54.3 78.6
Visionary-R1 3B RL QA 69.4 24.7 66.5 84.1
@@ 4.2 Main Results
The results are shown in Tab. 1. Comparing SFT with the base model, we observe that the performance of SFT is worse on three out of four datasets, with the biggest performance decline reaching 12% on MathVision. These results suggest that the model learned with question-answer pairs overfits the training data distribution. GRPO slightly outperforms the base model, achieving improvements of 0.3% on MathVista, 1.2% on MathVision, and 1.9% on MMStar. However, GRPO underperforms the base model by 1.5% on MMBench, which suggests that visual reasoning is difficult to learn from just question-answer pairs. By digging into the outputs, we observe that GRPO often leads to shortcuts in easy training samples while produces short, useless reasoning answers for unseen samples, as illustrated in Fig. 1.
Compared to SFT and GRPO, Visionary-R1 demonstrates huge potential in learning general-purpose reasoning capabilities, evidenced by the improvements of 7.9% on MathVista, 5.6% on MathVision, 14.1% on MMStar, and 2% on MMBench, over the base model. Compared with reasoning models that rely on labeled reasoning data, Visionary-R1 still maintains clear advantages on most datasets, despite using only question-answer pairs. Notably, Visionary-R1 even surpasses strong commercial AI models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, on MathVista, MMStar, and MMBench. These results strongly justify the effectiveness of learning to caption before reasoning.

### v3

@@ Abstract
Learning general-purpose reasoning capabilities has long been a challenging problem in AI. Recent research in large language models (LLMs), such as DeepSeek-R1, has shown that reinforcement learning techniques like GRPO can enable pre-trained LLMs to develop reasoning capabilities using simple question-answer pairs. In this paper, we aim to train visual language models (VLMs) to perform reasoning on image data through reinforcement learning and visual question-answer pairs, without any explicit chain-of-thought (CoT) supervision. Our findings indicate that simply applying reinforcement learning to a VLM—by prompting the model to produce a reasoning chain before providing an answer—can lead the model to develop shortcuts from easy questions, thereby reducing its ability to generalize across unseen data distributions. We argue that the key to mitigating shortcut learning is to encourage the model to interpret images prior to reasoning. Therefore, we train the model to adhere to a caption-reason-answer output format: initially generating a detailed caption for an image, followed by constructing an extensive reasoning chain. When trained on 273K CoT-free visual question-answer pairs and using only reinforcement learning, our model, named Visionary-R1, outperforms strong multimodal models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, on multiple visual reasoning benchmarks. Code and models will be publicly released.
@@ 1 Introduction
Reasoning is essential for enabling AI to tackle complex problems and make informed decisions in real-world applications. However, training AI models to reason is extremely challenging—primarily due to the lack of large-scale human-annotated reasoning data (Lightman et al., 2023; Christiano et al., 2017; Ouyang et al., 2022). Recent advances in large language models (LLMs), such as DeepSeek-R1 (Guo et al., 2025a), have demonstrated the potential to induce reasoning capabilities in LLMs via reinforcement learning and using only question-answer pairs, without explicit step-by-step supervision. Meanwhile, the computer vision community has begun exploring RL approaches for visual language models (VLMs), using methods like GRPO (Shao et al., 2024) to extend reasoning to multimodal settings (Meng et al., 2025; Feng et al., 2025; Liu et al., 2025; Shen et al., 2025). While these efforts are promising, existing visual reasoning models often rely on complex multi-stage training pipelines that are both computationally expensive and time-consuming. Moreover, these models heavily rely on labeled chain-of-thought reasoning data distilled from proprietary models like GPT-4o—limiting scalability and openness.
@@
To address the shortcut issue, we propose Visionary-R1, a reinforcement learning framework that enforces visual understanding before reasoning. The key idea is to train the model in a structured caption–reason–answer format, where it must first generate a detailed caption of the image before reasoning and answering. The captioning step ensures that the model does not just rely on superficial cues or patterns but engages in a deeper analysis of the image context, regardless of whether the question is easy or hard—this forces the model to adopt a consistent problem-solving approach, thus mitigating potential shortcuts and consequently making the reasoning capabilities more generalizable across different data distributions. To ensure the caption is informative, we impose auxiliary supervision on the caption tokens by using reinforcement learning from AI feedback (Bai et al., 2022). This caption reward is combined with standard accuracy and format rewards during policy optimization. The resulting model produces longer, more meaningful reasoning tokens than the model learned with GRPO alone (see Fig. 1), leading to better generalization performance on unseen data (see Fig. 2).
To evaluate our approach, we compile a comprehensive dataset that aggregates 11 popular question-answer datasets, covering areas such as scene understanding, chart analysis, mathematical problem-solving, and document processing. In total, the training data consists of 272.6K CoT-free question-answer pairs. After training, Visionary-R1 is evaluated on several challenging visual reasoning benchmarks including MathVista (Lu et al., 2023), MathVision (Wang et al., 2024), MMBench (Liu et al., 2024), MMMUPro (Yue et al., 2024), MMStar (Chen et al., 2024a), and CV-Bench (Tong et al., 2024). The results show that Visionary-R1 outperforms strong proprietary models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, as well as the latest competitors based on supervised pre-training and reinforcement fine-tuning.
@@ 2 Related Work
Learning LLMs/VLMs that can reason have gained increasing attention from both academia and industry due to their ability to generate human-like, step-by-step reasoning, which is advantageous for tackling complex problems and delivering more interpretable answers (Wei et al., 2022; Kojima et al., 2022). Supervised fine-tuning (SFT) is the most straightforward method to enhance a model’s reasoning capabilities, which relies on labeled data containing thinking processes. Since collecting human annotations is costly, existing work often resorts to using a pre-trained model like OpenAI’s GPT-4o to generate reasoning labels. For instance, LLaVA-CoT (Xu et al., 2024) utilizes GPT-4o to label 100K visual question-answer datasets with detailed chain-of-thought including summary, caption, and reasoning. However, the process of collecting CoT labels can be quite expensive, and the use of GPT-4o limits scalability while introducing a significant performance upper bound. Similarly, MMCR (Yan et al., 2025) also creates a 310k multi-turn reasoning dataset using GPT-4o. CoMCTS (Yao et al., 2024a) introduces the Mulberry-260k dataset, which is specifically crafted to train tree-structure reasoning models. Compared to these models, our Visionary-R1 only uses simple question-answer pairs for training without any chain-of-thought supervision, yet it achieves stronger reasoning performance.
@@
Compared to SFT, reinforcement learning (RL) has recently been proved more effective in developing general-purpose reasoning capabilities as this paradigm has the potential to enable the model to explore reasoning in a broader language space and develop its own thinking processes (Chu et al., 2025). Insight-V (Dong et al., 2024) presents a multi-agent system to select preference data from self-generated reasoning paths and optimizes the model based on a preference learning algorithm. R1-VL (Zhang et al., 2025) designs step-wise rewards to improve reasoning accuracy and validity but relies on labeled data for SFT. RL has also been applied in Vision-R1 (Huang et al., 2025) and R1-Onevision (Yang et al., 2025), but only 10K samples are used in these models for RL training while the main focus is on SFT (that uses more than 200K samples). Similarly, the Pixel Reasoner (Su et al., 2025) and VL-Rethinker (Wang et al., 2025) encourage deeper reasoning through images or explicit textual self-reflection, but their training pipelines still heavily rely on SFT with complex dataset selection and annotation processes. Our Visionary-R1 departs from the popular SFT-followed-by-RL pipeline and adopts a pure RL approach, eliminating reliance on large-scale annotated datasets required for SFT and enables more flexible, autonomous reasoning through RL-from-AI-feedback.
@@ 3.1 Motivation: The Shortcut Phenomenon in Visual Reasoning
While the GRPO (Shao et al., 2024) algorithm has been shown effective in improving the reasoning capabilities of language models, we observe a critical failure mode when transferring to visual reasoning tasks. This phenomenon manifests as a shortcut—GRPO often leads to degenerate behaviors where the model ignores the visual input and relies primarily on textual patterns from the question to generate an answer. As shown in Fig. 1, the model trained with GRPO can produce correct answers for simple questions during training—yet this is achieved without grounding in the image. This shortcut behavior can be particularly problematic in visual reasoning tasks, where the correct answer often depends on subtle image features such as embedded text, numerical values, object relationships, or chart patterns. Without forcing the model to attend to these visual signals, reinforcement learning alone encourages reward hacking: the model learns to exploit training distribution artifacts instead of learning general-purpose reasoning. To address this challenge, we propose a simple but effective modification: force the model to explicitly interpret the image before it begins reasoning. We operationalize this through a caption reward design (Section 3.2), which is then explicitly incorporated into the RL training objective (Section 3.3).
@@ 3.2 Visionary-R1: Grounding Reasoning via Captioning
While the format enforces structure, it does not guarantee that the caption is sufficiently detailed to support reasoning. To address this issue, we introduce a specialized caption reward $r_{c}\in\{0,1\}$ based on reinforcement learning from AI feedback (Bai et al., 2022). Specifically, we feed the generated caption into an LLM, and ask it to answer the question based solely on the caption. In implementation, we use the LLM component of the policy model. If the answer is correct, the caption is deemed informative and rewarded; otherwise, it is penalized. This encourages the model to produce useful, visually grounded descriptions. The final reward for a sampled sequence $i$ is computed as:
@@ 3.3 Training Objective with Caption Reward
Group Relative Policy Optimization, known as GRPO, was originally developed in DeepSeekMath (Shao et al., 2024) for text-only reasoning tasks, and later adopted in DeepSeek-R1 (Guo et al., 2025a). GRPO simplifies the reinforcement learning paradigm by getting rid of the critic model. This is done by generating a group of responses for each sample and then computing the normalized reward within the group to determine an advantage value. To adapt it to visual reasoning, our method introduces two key differences. (1) First, as described in Section 3.2, we design a new reward structure by adding a caption reward that explicitly evaluates whether the model has interpreted the visual input, addressing the shortcut issue. (2) Second, we introduce a cosine-annealed KL penalty to stabilize training and encourage longer, more meaningful outputs—avoiding the limitations of a static KL coefficient in multimodal settings. We now detail our training objective and implementation.
@@
The KL divergence in Eq. 4 serves as a penalty term to prevent the model from straying too far from the baseline policy model, thereby stabilizing the training. It is non-trivial to determine the balancing weight for this term: using a large weight forces the model to stay within a close neighborhood of the baseline model and therefore impedes the model’s ability to engage in more in-depth thinking and generating long, detailed reasoning; on the other hand, using a small weight can lead to unstable training and potentially result in reward hacking (Skalse et al., 2022). To overcome this challenge, we propose dynamically annealing the KL penalty coefficient over time using cosine annealing, which uses a large coefficient during the early, unstable training phase and gradually reduces the value to allow the model to produce longer outputs in later stages. Specifically, we replace $\beta$ in Eq. 3 with $\hat{\beta}$, which is calculated as
@@ 4.1 Experimental Settings
We evaluate our approach on several widely-used visual reasoning benchmarks that cover various visual formats and question types: MathVista (Testmini) (Lu et al., 2023), MathVision (Wang et al., 2024), and MMBench (en) (Liu et al., 2024). MathVista encompasses a variety of reasoning types, including logical, algebraic, and scientific reasoning questions. MathVision focuses on mathematical visual reasoning tasks. MMBench is a comprehensive evaluation suite concerned with visual and mathematical reasoning. Meanwhile, we also included the results of MMMUPro (Yue et al., 2024), MMStar (Chen et al., 2024a) and CV-Bench (Tong et al., 2024) in the  A.4 to provide a more diverse and comprehensive evaluation.
@@
We adopt Qwen2.5-VL-3B (Bai et al., 2025) as the base model. This pre-trained model has strong visual understanding capabilities but has not undergone post-training for reasoning. For the group reward computation, we generate 8 output sequences (i.e., $n=8$ in Eq. 3) and the sampling temperature is set to 0.9 following the common practice. All parameters are optimized with a learning rate of $5\times 10^{-7}$. The caption reward’s balancing weight $\alpha$ is set to 0.1. The KL coefficient $\beta$ is set to 0.04.
Table 1: Comparison with state-of-the-arts on three challenging visual reasoning benchmarks. SFT and RL mean supervised fine-tuning and reinforcement learning, respectively. CoT means chain-of-thought, which is either self-generated or distilled from third-party models like GPT-4o. QA means that the model is learned with question-answer pairs only. Despite having only 3B parameters and using only QA data for training, Visionary-R1 beats strong commercial models like GPT-4o and Claude3.5-Sonnet. Note that indicates results borrowed from the Seed’s report (Guo et al., 2025b).
Size Strategy Data MathVista MathVision MMBench
@@
GPT-4o (Hurst et al., 2024) - - - 63.8 31.2 84.3
GPT-o1 (Jaech et al., 2024) - - - 71.8 63.2 83.8
Claude3.5-Sonnet (Anthropic, 2024) - - - 67.7 37.9 82.6
Claude3.7-Sonnet (Anthropic, 2025) - - - 74.5 58.6 82.0
Gemini-1.5-Pro (Team et al., 2024) - - - 63.9 19.2 73.9
Gemini-2.5-Pro (Google, 2025) - - - 82.7 73.3 90.1
@@
Qwen2.5-VL (Bai et al., 2025) 3B - - 62.3 21.2 79.1
InternVL2.5 (Chen et al., 2024b) 4B - - 60.5 20.9 81.1
MiniCPM-V2.6 (Yao et al., 2024b) 8B - - 60.6 17.5 81.5
LLaMA3.2 (AI, 2024) 11B - - 51.5 - 65.8
@@
Ovis (Yan et al., 2025) 4B SFT CoT 66.6 - 79.3
Mulberry (Yao et al., 2024a) 7B SFT CoT 63.1 - -
R1-Onevision (Yang et al., 2025) 7B SFT+RL CoT 64.1 29.9 -
Insight-V (Dong et al., 2024) 7B SFT+RL CoT 59.9 - 82.3
R1-VL (Zhang et al., 2025) 7B SFT+RL CoT 63.5 24.7 -
LLaVA-CoT (Xu et al., 2024) 11B SFT CoT 54.8 - 75
@@
Base Model 3B - - 61.5 19.1 82.1
SFT 3B SFT QA 54.6 7.0 80.7
GRPO 3B RL QA 61.8 20.3 78.6
[…]
@@ 4.2 Main Results
The results are shown in Tab. 1. Comparing SFT with the base model, we observe that the performance of SFT is worse on three out of four datasets, with the biggest performance decline reaching 12% on MathVision. These results suggest that the model learned with question-answer pairs overfits the training data distribution. GRPO slightly outperforms the base model, achieving improvements of 0.3% on MathVista, 1.2% on MathVision. However, GRPO underperforms the base model by 1.5% on MMBench, which suggests that visual reasoning is difficult to learn from just question-answer pairs. By digging into the outputs, we observe that GRPO often leads to shortcuts in easy training samples while produces short, useless reasoning answers for unseen samples, as illustrated in Fig. 1.
Compared to SFT and GRPO, Visionary-R1 demonstrates huge potential in learning general-purpose reasoning capabilities, evidenced by the improvements of 7.9% on MathVista, 5.6% on MathVision, and 2% on MMBench, over the base model. Compared with reasoning models that rely on labeled reasoning data, Visionary-R1 still maintains clear advantages on most datasets, despite using only question-answer pairs. Notably, Visionary-R1 even surpasses strong commercial AI models, such as GPT-4o, Claude3.5-Sonnet, and Gemini-1.5-Pro, on MathVista, and MMBench. These results strongly justify the effectiveness of learning to caption before reasoning.